# Topological Phases, Criticality, and Mixed State Order in a Hubbard Quantum Simulator

### v1

@@ Abstract
Phases of matter are conventionally distinguished from one another by local observables. Topological quantum phases lie outside this paradigm; their differences can only be learned by examining them globally. This has striking implications for the stability of these phases, their classification, and the phase transitions between them. In this work, we experimentally demonstrate these implications using interacting magnetic erbium atoms in an optical lattice [1, 2]. We show that a Mott insulator and a pinned charge-density wave in one dimension are in distinct crystalline symmetry-protected topological phases (CSPTs) [3, 4, 5, 6, 7]. The quantum phase transition separating them is revealed by measuring nonlocal string order parameters using site-resolved imaging. Remarkably, stacking two copies of these states eliminates the critical point—a signature feature of topological phases that underlies their classification [8, 9, 10, 11, 12, 13]. Moreover, we show that while programmable symmetry-breaking disorder can also remove this critical point, averaging over disorder restores it, supporting recent theoretical predictions of mixed-state order [14, 15, 16, 17, 18, 19]. Finally, we highlight a connection between one of these CSPTs and the Haldane insulator [20, 21], and detect signatures of the transition between the Haldane and the Mott insulator. Our results establish a path toward probing broader symmetry-protected topology and mixed-state order in programmable quantum devices.
@@ I Introduction
Figure 1: Topological phases and criticality in a staggered dipolar Bose-Hubbard quantum simulator. a, With long-range density-density interactions ($V$), a symmetry protected topological (SPT) phase, named the Haldane insulator (HI), emerges in the 1D extended Bose-Hubbard phase diagram [21]. The HI is in a distinct phase from the Mott insulator (MI). Remarkably, the topological transition persists when introducing a staggered chemical potential $\mu_{S}$ [6]. In this regime, the HI and MI admit atomic limits which are in two distinct crystalline SPTs (CSPTs) protected by the parity conservation of bosons Eq. (2)) and site-centered inversion symmetry (dashed line). b, Experimentally, we realize this model using bosonic 164Er atoms in an optical lattice. Particles tunnel with amplitude $t$ and experience on-site interactions of strength $U$. Dipolar interactions between erbium atoms induce long-range density-density coupling with nearest-neighbor strength $V$. A superlattice generates a staggered chemical potential $\mu_{S}$. We directly resolve the atom number per site without parity projection. A representative single-shot image is shown, with the digitized filling below. The camera-count histogram displays peaks at 00, $1$, and $2$ particles per site. c, We probe the stability of the topological criticality by stacking two instances of the 1D chain, demonstrating that these quantum phases of matter are invertible [22]. d, Moreover, we demonstrate that the criticality is unstable to programmable symmetry-breaking disorder, but is restored when averaging over disorder. In this sense, these SPTs correspond to distinct mixed-state quantum phases of matter [18].
Understanding the quantum phases of matter that can emerge from strongly correlated quantum systems is one of the most compelling and active frontiers of research in contemporary physics. A key avenue of investigation has centered around analyzing the phase diagrams of the Bose and Fermi-Hubbard models as they provide a minimal framework for capturing the interplay between kinetics and strong inter-particle interactions. In this context, analog quantum simulators—based on ultracold atoms in optical lattices—are emerging as invaluable tools [23]. Indeed, these platforms provide natural realizations of such models and enable single-site resolved measurements [24, 25]. Consequently, a plethora of exotic phases have been seen [26, 27].
Many of these phases fall within Landau’s paradigm of spontaneous symmetry breaking [28, 29]: they are equilibrium states that fail to respect a symmetry of their Hamiltonian and are diagnosed via local observables [30]. In contrast, we focus on one-dimensional bosonic symmetry-protected topological phases (SPTs) [31, 32, 10, 11, 12, 13], in particular, the Haldane insulator [20, 21] and crystalline SPTs (CSPTs) shown in Fig. 1a. These are interacting phases that do not break any symmetries and cannot be distinguished by local correlation functions. Such phases are instead only distinguished by measuring non-local correlations present within the system and are distinct from other phases provided that their protecting symmetries remain unbroken.
Rapid progress has been made in the experimental exploration of SPTs. In particular, several pioneering solid-state experiments indirectly identified such states by locally probing the decoupled edge state present in some 1D SPTs [33, 34, 35, 36, 37, 38]. More recently, advances in quantum simulation—e.g., in ultracold atoms [39, 40, 41, 42], superconducting circuits [43, 44, 45, 46], and trapped ions [47]— have enabled a characterization of their non-local correlations and hence their SPT order.
Beyond their identification, several frontiers in our understanding of SPTs remain to be explored experimentally, specifically regarding the properties of the topological phase transitions between them and the stability of these states to symmetry-breaking disorder. In particular, unlike conventional phase transitions, theory predicts that the topological transition between SPTs (and consequently the features that distinguish them) can be removed by stacking several copies of the system—a fact that lies at the heart of the classification of SPTs [32, 8, 9, 10, 11, 12, 13] (Fig. 1c). Moreover, while SPTs are typically only stable in the presence of their protecting symmetries, in certain cases, they can remain stable even in the presence of symmetry-breaking disorder, provided that the symmetry is present upon disorder averaging [14, 15, 16, 17, 18, 19, 48] (Fig. 1d). Since the ensemble of pure quantum ground states for each disorder pattern defines a mixed state, this implies that certain SPTs are mixed state quantum phases [49, 50, 51, 52, 53, 54, 55, 56], which have recently garnered significant theoretical interest but have yet to be studied experimentally. Here, we show that by probing a variant of the Bose–Hubbard model, we can identify topological transitions, remove the criticality by stacking, and probe mixed state orders.
We consider a Bose-Hubbard model with long-range interactions and a staggered chemical potential:
$\begin{split}H=-t\sum_{x}(\hat{a}_{x}^{\dagger}\hat{a}_{x+1}+\mathrm{h.c.})+\frac{U}{2}\sum_{x}\hat{n}_{x}(\hat{n}_{x}-1)\\ +\sum_{x,r>0}V_{r}\hat{n}_{x}\hat{n}_{x+r}-\mu_{S}\sum_{x}(-1)^{x}\hat{n}_{x},\end{split}$ (1)
@@
where $t$ and $U$ denote the tunneling and on-site interaction energies, respectively, $\mu_{s}$ is the strength of a staggered potential, $a_{x}^{\dagger}$ ($a_{x}$) is the bosonic creation (annihilation) operator at site $x$, and $\hat{n}_{x}=a_{x}^{\dagger}a_{x}$ is the bosonic number operator. $V_{r}$ describes the long-range interaction between particles separated by a distance $r$ (Methods). We refer to this model as the staggered dipolar Bose-Hubbard model.
@@
To simulate this Hamiltonian, we use magnetic erbium atoms with dipolar interactions, and tune the on-site interaction via magnetic Fano-Feshbach resonances [57] (Fig. 1b). To access low-temperature 1D states with a high data acquisition rate, we adiabatically load bosonic 164Er atoms into 2D optical lattices with a large lattice depth in one direction, effectively realizing several decoupled 1D systems (Methods). Thousands of experimental realizations are performed, followed by data selection procedures (Methods) that include restricting the analysis to samples with unit filling. The boundary conditions are described in Methods. We then extract observables from site-resolved images that directly measure atom number per site without parity projection [58].
@@ II Crystalline SPTs and Their Transitions
Figure 2: Quantum phase transition between two crystalline symmetry protected topological phases (CSPTs). We investigate the quantum phase transition between two CSPTs in one-dimensional chains of 16 sites. a, For the atomic limit CSPT, the parity of an odd-length string for a unity-filled state has $-1$ parity. Even if tunneling takes us away from this atomic-limit state, the parity remains negative in the presence of site-centered inversion symmetry. b, The sign of the parity string of odd length distinguishes between the two CSPTs and can be used to form effective “string order parameters” ($p_{1}$ and $p_{2}$). As $\mu_{S}$ crosses the quantum critical point, the system transitions from the CSPT-1 phase (Mott insulator-like, bottom) to the CSPT-2 phase (Charge Density Wave-like, top). Our experimentally measured order parameters show good agreement with the results from simulations taking into account readout infidelities (Methods). c, The experimental observables are plotted with error bars (bootstrapping standard error) and compared to simulations (shades indicate 95% confidence interval in this entire paper). On the CSPT-1 side of the transition, the $p_{2}$ observable (orange) is effectively zero (limited by finite imaging fidelity), while the local average density observable $n_{2}$ (brown) is significantly non-zero, highlighting the topological nature of the CSPTs. This shows that local observables cannot differentiate between the two CSPTs, whereas the parity “string order parameter” can. The non-sharpness of the local observables is not due to finite-size effects (Methods).
We begin our exploration of SPTs and their topological phase transitions by highlighting two symmetries of the Hamiltonian that are responsible for protecting the topological phases we study. First, as a consequence of particle number conservation, the system preserves the total parity (even or oddness) of the number of bosons in the chain. This yields a $\mathbb{Z}_{2}$ symmetry of the Hamiltonian generated by:
@@
With symmetries in hand, the topological phases we study can now be understood. These phases arise in the phase diagram of our system at unit filling $\bar{n}=1$ and with a large on-site Hubbard interaction compared to tunneling, as shown in the upper section of the phase diagram in Fig. 1a. The first CSPT phase (CSPT-$1$) is essentially the Mott insulator, which appears for small values of the staggered chemical potential. Here, the strong on-site Hubbard interaction stabilizes an average number of bosons on every site close to $1$. In contrast, the second CSPT (CSPT-$2$) is found when the staggered chemical potential is large, and consequently, there is an average of two bosons on the odd sites (‘doublons’) and no bosons on the even sites. These configurations are schematically depicted in the phase diagram of Fig. 1a, where the classical states (|$\ket{1111\cdots}$ and |$\ket{2020\cdots}$) become exact in the atomic limit, i.e., when the tunneling $t\to 0$.
@@
These two states cannot be adiabatically connected in the presence of the aforementioned parity $\hat{\mathcal{P}}$ and inversion $\hat{\mathcal{I}}$ [6]. While there exists no local order parameter to detect these two phases in the presence of quantum fluctuations [6], we show that they are instead distinguished by a nonlocal (‘topological’) order parameter. In Sec. V we discuss how CSPT-2 is adiabatically connected to the Haldane insulator, as already hinted at in Fig. 1a.
To understand this invariant, let us first work with the product states that appear in the atomic limit. One way to distinguish these states is to recognize that in any large inversion-symmetric region, the parities of bosons in CSPT-1 and 2 are $-1$ (odd) and $+1$ (even) respectively (Fig. 2a for CSPT-1, top). As quantum fluctations are introduced and we move away from the atomic limit, the bosons in these states delocalize into dressed quasiparticles (Fig. 2a, bottom) and the expectation value of the parity in this region is no longer exactly $+1$ or $-1$. Nevertheless, the sign of the expectation value of the parity in these regions is a discrete invariant that can be used to distinguish the two. As long as the gap stays open and both parity and inversion symmetry are preserved, any change in the parity expectation within the region occurs symmetrically from its edges. As a result, the sign of the bulk parity remains unchanged (Fig. 2a, bottom).
@@
Figure 3: Elimination of the phase transition via interchain coupling. Coupling an even number of $\mathbb{Z}_{2}$ topological chains can trivialize the topology. a, To quantum simulate the pair-creation coupling in Eq. (4), we map our model (top row) to a half-filling representation (middle row), and then apply a particle-hole transformation to the bottom chain (bottom row). Our system size is 2 by 10 sites. b, When inter-chain tunneling $t_{\perp}$ is nonzero, the numerically computed energy gap at the critical point of a single CSPT chain (e.g., $\mu_{S}/t=0$) becomes finite. To probe this gap, we initialize two CSPT-1 chains at minimal $t_{\perp}$ and ramp into the CSPT-2 regime via various intermediate values. We vary the staggered chemical potential $\mu_{S}$ linearly in time while simultaneously ramping $t_{\perp}$ to a final configuration with negligible coupling, and read out snapshots of the resulting state. c, At the final $\mu_{S}=6t$ point, tunneling is negligible, and the density of states can be approximated classically. We analyze each snapshot to compute energy histograms of the final state. As $t_{\perp}$ increases, the extracted temperature decreases. Colored bars show data; colored dotted lines are Gibbs distribution fits; the gray dotted line shows the infinite-temperature distribution. d, Measured temperatures (error bars: central 68% percentile via bootstrap) systematically decrease with increasing $t_{\perp}$. e, Parity order parameters versus the maximum $t_{\perp}$ (error bars: bootstrapped standard error) reveal a crossover into the CSPT-2 phase only at sufficiently large coupling, consistent with the presence of a gapped path.
@@
To demonstrate the two different topological phases in our quantum simulator, we experimentally measure [59, 60] this topological invariant via the “string operator” defined as:
@@
which measures the average parity of bosons in a region. The experimental results are shown in Fig. 2b as a function of the staggered field over the interaction strength, $\mu_{S}/U$. As expected, when the staggered field is small, the parity is negative for odd-length strings (encoding the topological invariant) and is positive for even-length strings. In contrast, when the staggered field is large, we observe that the parity is positive for all length strings. These experimental results agree with numerically simulated results (Methods).
The sharp distinction between these phases can be diagnosed by computing the staggered and uniform parity “string order parameters” $p_{1}$ and $p_{2}$, defined by averaging $(-1)^{r}\mathcal{P}_{x_{0}}(r)$ and $\mathcal{P}_{x_{0}}(r)$ respectively over sufficiently long $r$ (Methods). These order parameters are shown in the top panel of Fig. 2c. We observe that they clearly distinguish between two phases—they are either finite or effectively zero—with a sharp transition in between at the critical point. One way to argue that these order parameters vanish outside of their respective phases is using an approximate particle-hole symmetry of the model in this triplon-free regime (Methods). In contrast, we also define local observables of the average density fluctuations defined as $n_{2}=\langle(-1)^{i}\delta\hat{n}_{i}\rangle$, where $\delta\hat{n}_{i}=\hat{n}_{i}-1$, and $n_{1}=1-n_{2}$ (Fig. 2c bottom). Their smooth variation across the phase transition reflects the inability of local observables to faithfully distinguish between the two phases — a defining feature of topological phases (Methods). These observations strongly support the existence of two distinct CSPTs at $\mu_{S}\ll U$ and $\mu_{S}\gg U$, separated by a quantum phase transition, in agreement with our large-scale ground state numerical simulations [61] (Methods).

### v2

# Topological Phase Transitions and Mixed State Order in a Hubbard Quantum Simulator

@@ Abstract
Topological phase transitions challenge conventional paradigms in many-body physics by separating phases that are locally indistinguishable yet globally distinct. Using a quantum simulator of interacting erbium atoms in an optical lattice, we observe such a transition between one-dimensional crystalline symmetry-protected topological phases (CSPTs). We detect the critical point through non-local string order parameters and reveal its connection to the transition predicted between the Mott and Haldane insulators. Moreover, we demonstrate a striking property: stacking two identical systems eliminates the transition, confirming the predicted group structure and invertibility of SPTs. Finally, while introducing symmetry-breaking disorder also removes the transition, disorder averaging restores it. Consequently, the adjacent phases realize a form of mixed-state quantum order wherein the criticality between them depends on the observer’s information. Our results demonstrate how topology and information influence quantum phase transitions, opening the doors to probing novel critical phenomena in programmable quantum matter.
@@ I Introduction
Figure 1: Topological phases and criticality in a staggered dipolar Bose-Hubbard quantum simulator. a, With long-range density-density interactions ($V$), a symmetry protected topological (SPT) phase, named the Haldane insulator (HI), emerges in the 1D extended Bose-Hubbard phase diagram [1, 2]. The HI is in a distinct phase from the Mott insulator (MI). Remarkably, the topological transition persists when a staggered chemical potential $\mu_{S}$ is introduced. In this regime, the HI and MI admit atomic limits which are in two distinct crystalline SPTs (CSPTs) protected by the parity conservation of bosons (Eq. (2)) and site-centered inversion symmetry (dashed line) [3]. b, Experimentally, we realize this model using bosonic 164Er atoms in an optical lattice. Particles tunnel with amplitude $t$ and experience on-site interactions of strength $U$. Dipolar interactions between erbium atoms induce long-range density-density coupling with nearest-neighbor strength $V$. A superlattice generates a staggered chemical potential $\mu_{S}$. We directly resolve the atom number per site without parity projection. A representative single-shot image is shown, with the digitized filling below. The camera-count histogram displays peaks at 00, $1$, and $2$ particles per site. c, We probe the stability of the topological criticality by stacking two instances of the 1D chain, demonstrating that these quantum phases of matter are invertible. d, Moreover, we demonstrate that the criticality is unstable to programmable symmetry-breaking disorder, but is restored when averaging over disorder. In this sense, these SPTs correspond to distinct mixed-state quantum phases of matter [4].
Systems of many interacting particles in equilibrium can undergo phase transitions where macroscopic properties abruptly change [5]. Traditionally, these transitions are described by the Landau paradigm, which posits that transitions occur when the correlations of local observables (or “order parameters”) become long-range and spontaneously break the system’s symmetries. This theory successfully explains many finite-temperature transitions, such as between liquids and crystals, or between different types of magnets. However, at near-zero temperatures, quantum effects lead to “topological” phase transitions with no temperature-driven analog, where nonlocal quantum correlations separate phases that are indistinguishable by local measurements.
Remarkably, such topological transitions have been predicted even for one-dimensional quantum systems, namely between symmetry-protected topological (SPT) phases [6, 7, 8, 9, 10, 11, 12, 13]. Recent advances in analog quantum simulation have enabled the realization of such 1D SPT phases—in platforms ranging from ultracold atoms [14, 15, 16, 17], superconducting circuits [18], and trapped ions [19]—in addition to digital simulations [20, 21, 22]. However, detecting the quantum critical points between these phases has remained an outstanding challenge in analog simulators, requiring a tunable parameter to sweep across the transition.
In this work, we realize such quantum critical points between SPT phases in an analog quantum simulator (Fig. 1a) and demonstrate their topological nature by probing their stability and instability under stacking and disorder (Fig. 1c, d). We employ a quantum simulator [23, 24, 25] to realize a variant of the Bose–Hubbard model featuring long-range interactions and a staggered chemical potential (Fig. 1b):
$\begin{split}H=-t\sum_{x}(\hat{a}_{x}^{\dagger}\hat{a}_{x+1}+\mathrm{h.c.})+\frac{U}{2}\sum_{x}\hat{n}_{x}(\hat{n}_{x}-1)\\ +\sum_{x,r>0}V_{r}\hat{n}_{x}\hat{n}_{x+r}-\mu_{S}\sum_{x}(-1)^{x}\hat{n}_{x}.\end{split}$ (1)
@@
Here, $t$, $U$, and $\mu_{S}$ denote the tunneling, on-site interaction, and staggered potential strength, respectively; $a_{x}^{\dagger}$ ($a_{x}$) creates (annihilates) a boson at site $x$, and $\hat{n}_{x}=a_{x}^{\dagger}a_{x}$ is the number operator. $V_{r}$ describes long-range dipolar interactions at distance $r$ (see Supplementary Materials, SM). We refer to this model as the staggered dipolar Bose-Hubbard model. Previous work [1, 2] has predicted a Haldane insulator [26] in its phase diagram, which we also probe.
@@
To simulate this Hamiltonian, we use magnetic erbium atoms with dipolar interactions [27] and tune the on-site interaction via Fano-Feshbach resonances [28] (Fig. 1b). We adiabatically load bosonic 164Er atoms into 2D optical lattices with deep confinement in one direction to create isolated 1D chains (SM). Thousands of experimental realizations are performed, followed by data selection procedures (SM) that include restricting the analysis to samples with unit filling. The boundary conditions are described in the SM. Site-resolved imaging [29, 30] then yields atom-number measurements without parity projection [31].
@@ II Crystalline SPTs and Their Transitions
Figure 2: Quantum phase transition between crystalline symmetry protected topological phases. We investigate the quantum phase transition between two CSPTs in one-dimensional chains of 16 sites. a, For the atomic limit CSPT, the parity of an odd-length string for a unity-filled state has $-1$ parity. Even if quantum fluctuations, e.g., induced by tunneling, take us away from this atomic-limit state and the bosons mildly delocalize, the parity remains negative in the presence of site-centered inversion symmetry. b, The sign of the parity string of odd length distinguishes between the two CSPTs and provides a topological invariant. As $\mu_{S}$ crosses the quantum critical point, the system transitions from the CSPT-1 phase (Mott insulator-like, bottom) to the CSPT-2 phase (Charge Density Wave-like, top). Our experimentally measured parity string shows good agreement with the results from simulations taking into account readout infidelities (SM). c, We use the parity string measurements to define two effective “string order parameters” for CSPT-1 and 2 ($p_{1}$ and $p_{2}$) shown in the top panel. The experimental observables are plotted with error bars (bootstrapping standard error) and compared to simulations (shades indicate 95% confidence interval in this entire paper). On the CSPT-1 side of the transition, the $p_{2}$ observable (orange) is effectively zero (limited by finite imaging fidelity), while the local average density observable $n_{2}$ (brown) is significantly non-zero, highlighting the topological nature of the CSPTs. This shows that local observables cannot differentiate between the two CSPTs, whereas the parity “string order parameter” can. The crossover of the local observables is not due to finite-size effects (SM).
We begin by highlighting two symmetries protecting the SPT phases and their transitions. First, due to particle number conservation, the system preserves the total parity (even or oddness) of the number of bosons in the chain. This yields a $\mathbb{Z}_{2}$ symmetry of the Hamiltonian generated by:
@@
These phases arise in the phase diagram of our system at unit filling $\bar{n}=1$ and with a large on-site Hubbard interaction compared to tunneling, as shown in the upper section of the phase diagram in Fig. 1a. The first CSPT phase (CSPT-$1$) is essentially the Mott insulator, which appears for small values of the staggered chemical potential. Here, the strong on-site Hubbard interaction stabilizes an average number of bosons on every site close to $1$. In contrast, the second CSPT (CSPT-$2$) is found when the staggered chemical potential is large, and consequently, there is an average of two bosons on the odd sites (‘doublons’) and no bosons on the even sites. These configurations are schematically depicted in the phase diagram of Fig. 1a, where the classical states (|$\ket{1111\cdots}$ and |$\ket{2020\cdots}$) become exact in the atomic limit, i.e. when the tunneling $t\to 0$.
@@
These two states cannot be adiabatically connected in the presence of the aforementioned parity $\hat{\mathcal{P}}$ and inversion $\hat{\mathcal{I}}$ [3]. Indeed, while there exists no local order parameter to distinguish these two phases in the presence of quantum fluctuations [3], they are instead differentiated by a nonlocal topological invariant. This invariant is precisely the sign of the average boson parity in a large region of odd length. In particular, this sign is negative (positive) in CSPT-1 (CSPT-2), and cannot change without closing the energy gap (Fig. 2a and SM).
@@
Figure 3: Elimination of the phase transition via stacking. Coupling an even number of $\mathbb{Z}_{2}$ topological chains can trivialize the topology. a, To quantum simulate the pair-creation coupling in Eq. (4), we map our model (top row) to a half-filling representation (middle row), and then apply a particle-hole transformation to the bottom chain (bottom row). Our system size is 2 by 10 sites. b, When inter-chain tunneling $t_{\perp}$ is nonzero, the numerically computed energy gap at the critical point of a single CSPT chain (e.g., $\mu_{S}/t=0$) becomes finite. To probe this gap, we initialize two CSPT-1 chains at minimal $t_{\perp}$ and ramp into the CSPT-2 regime via various intermediate values. We vary the staggered chemical potential $\mu_{S}$ linearly in time while simultaneously ramping $t_{\perp}$ to a final configuration with negligible coupling, and read out snapshots of the resulting state. c, At the final $\mu_{S}=6t$ point, tunneling is negligible, and the density of states can be approximated classically. We analyze each snapshot to compute energy histograms of the final state. As $t_{\perp}$ increases, the extracted temperature decreases. Colored bars show data; colored dotted lines are Gibbs distribution fits; the gray dotted line shows the infinite-temperature distribution. d, Measured temperatures (error bars: central 68% percentile via bootstrap) systematically decrease with increasing $t_{\perp}$. e, Parity order parameters versus the maximum $t_{\perp}$ (error bars: bootstrapped standard error) reveal a crossover into the CSPT-2 phase only at sufficiently large coupling, consistent with the presence of a gapped path.
@@
To demonstrate the two different topological phases in our quantum simulator, we experimentally measure [36, 37] this topological invariant via the parity “string operator” defined as:
@@
which precisely measures the average parity of bosons in a region. The experimental results are shown in Fig. 2b as a function of the staggered field over the interaction strength, $\mu_{S}/U$. As expected, for weak staggered fields, the parity is negative for odd-length strings—reflecting the topological invariant—and positive for even-length strings. For strong staggered fields, the parity is positive for all string lengths. These observations are consistent with numerical simulations (SM).
The sharp distinction between these phases can be diagnosed by computing the staggered and uniform parity “string order parameters” $p_{1}$ and $p_{2}$, defined by averaging $(-1)^{r}\mathcal{P}_{x_{0}}(r)$ and $\mathcal{P}_{x_{0}}(r)$ respectively over sufficiently long $r$ (SM). These order parameters are shown in the top panel of Fig. 2c. They clearly distinguish the two phases, being either finite or vanishing, with a sharp transition at the critical point. One way to argue that these order parameters vanish outside of their respective phases is by using an approximate particle-hole symmetry of the model in this triplon-free regime (SM). In contrast, we also define local observables of the average density fluctuations defined as $n_{2}=\langle(-1)^{i}\delta\hat{n}_{i}\rangle$, where $\delta\hat{n}_{i}=\hat{n}_{i}-1$, and $n_{1}=1-n_{2}$ (Fig. 2c bottom). Their smooth variation across the phase transition reflects the inability of local observables to faithfully distinguish between the two phases — a defining feature of topological phases (SM). These observations strongly support the existence of two distinct CSPTs at $\mu_{S}\ll U$ and $\mu_{S}\gg U$, separated by a quantum phase transition, in agreement with our large-scale ground state numerical simulations [38] (SM).